A student is using the elimination method to solve the system of equations below. What is the best first step?

4x - 5y = 2
2x + y = -3

Found 3 solutions by Alan3354, stanbon, Fombitz:
Answer by Alan3354(69443)

A student is using the elimination method to solve the system of equations below. What is the best first step?

4x - 5y = 2
2x + y = -3
-----------------
Either multiply the 2nd eqn by 2 or by 5.
Makes no difference.

Answer by stanbon(75887)

4x - 5y = 2
2x + y = -3
------
You have a choice::
1) If you want to solve for "y", Multiply thru the bottom
equation by 2.
--
2) I you want to solve for "x", Multiply thru the bottom
equation by 5.
--------------
Cheers,
Stan H.
--------------

Answer by Fombitz(32388)

Not sure there's a "best" first step but here are two equivalent first steps.
Multiply the second equation by 5 to eliminate y.
or
Multiply the second equation by -2 to eliminate x.